OpenStudy (anonymous):

How do you find an equation of the line that passes through the point and has the indicated slope m? (2,4); m = -1 With this one, isn't the formula y-y1=m(x-x1)

OpenStudy (amistre64):

thats as good as any other form of the equation of a line

OpenStudy (anonymous):

Okay, but when I figure it, I can't seem to get it to come out to what you got

OpenStudy (anonymous):

Is (2, 4) = 2 = x1 and 4 = y1?

OpenStudy (anonymous):

-1 would = m correct?

OpenStudy (amistre64):

y-4 = -1(x-2) ; do you wanna shift it to the slope intercept now?

OpenStudy (anonymous):

Okay, I switched the y-4 and had -4-y, and just input 0 That is where my mistake was then...So you can have many different equations then for these types of problems?

OpenStudy (anonymous):

I just have to show an equation at this point...

OpenStudy (amistre64):

yes; you can have as many equations as there are ways to move everything around. All the forms of the equations do is move stuff without changing any values

OpenStudy (anonymous):

Wow, and I noticed that y-4=-1(x-2) cannot be simplified; therefore will be an equation correct?

OpenStudy (amistre64):

its like having a box of crayons; it doesnt matter how you arrange them in the box; you still have the same box of crayons

OpenStudy (anonymous):

Nice... Your always such a good help.

OpenStudy (anonymous):

How is the family?

OpenStudy (amistre64):

yes, the point slope form that is there is a valid form of a line equation :)

OpenStudy (anonymous):

You helped me before, personally. JaniSmit...remember

OpenStudy (amistre64):

if its been longer than 3 days, its merged into the past memories lol

OpenStudy (amistre64):

family is ok; daughter decided to sprain her ankle tho

OpenStudy (anonymous):

Nice...poor little thing...hope she recovers quickly

OpenStudy (amistre64):

:) yeah, shes bound to break other parts tho

OpenStudy (anonymous):

I told you my husband is military and I'm helping him with his math classes...this is the last one..thank God

OpenStudy (anonymous):

So when I have to find an equation of the line that passes through the given points such as (2,1) and (2,5) I would use the same formula right?

OpenStudy (amistre64):

slope formula; which is just telling you to subtract one point from another

OpenStudy (anonymous):

Or, m=y2-y/x2-x1

OpenStudy (anonymous):

Gotcha...awesome...i caught that one ..woohoo..

OpenStudy (amistre64):

yeah, that one... (x , y ) -(x1 , y1) --------- (x-x1, y-y1)

OpenStudy (anonymous):

Okay, so with (2,1) and (2,5) I would do... m= 5 -1 / 2 - 2 = 4 / 0 = 4?????

OpenStudy (amistre64):

(2,5) -(2,1) ------ 0,4; slope = y/x = 4/0 The ONLYtime you get this is when you have a vertical line.

OpenStudy (amistre64):

when you have a vertical line then the equation is simply: x = whatever your x value was in your points

OpenStudy (anonymous):

So it equals 0 then? As if a fraction 4/0=0

OpenStudy (amistre64):

not zero; undefined

OpenStudy (anonymous):

Undefined when x1=x2............

OpenStudy (amistre64):

0 slope is like driving on a flat road; its just flat, no elevation changes

OpenStudy (amistre64):

4/0 is like jumping off a cliff; you only go in one direction

OpenStudy (anonymous):

That slope could be going in any direction correct? Up and down or sideways right..

OpenStudy (anonymous):

Funny, I got you exactly..i understand

OpenStudy (anonymous):

I have another tough one that i'm lost with ..you game?

OpenStudy (amistre64):

sure :)

OpenStudy (anonymous):

This one says to find an equation of the line that has slope m and y-intercept be? for the equation m = -2; b= -1 I am unsure which formula to use really..

OpenStudy (anonymous):

Maybe the y-y1=m(x-x1)

OpenStudy (amistre64):

the one thats called slope intercept form :) since it gives a slope and an intercept

OpenStudy (amistre64):

y = mx +b

OpenStudy (amistre64):

y = -2x -1

OpenStudy (anonymous):

Darnit, you beat me..I found some formulas just now....it said exactly what you said..your awesome

OpenStudy (amistre64):

yeah :)

OpenStudy (anonymous):

Wow, so these are literally just chosing equations that do not need to be solved..

OpenStudy (amistre64):

well, the solution IS the line that is created by them

OpenStudy (amistre64):

which is why they are called 'line' equations :)

OpenStudy (amistre64):

one solution is called a point; all solutions are called a line

OpenStudy (anonymous):

So, If I am trying to find 2 different slopes that are parallel to one another, do I just create 2 different equations?

OpenStudy (anonymous):

I do not recall this kind of formula

OpenStudy (amistre64):

yep; but they should have the same slopes

OpenStudy (amistre64):

ax+by = c ax + by = d are parallel lines

OpenStudy (amistre64):

when c=d they are the same line

OpenStudy (anonymous):

Basically, if i had to find an equation for (-1,3) that is parallel to the line passing through points (-2, -3) and (2,5), would that formula fit that description?

OpenStudy (amistre64):

perpendicular lines are: ax + by = c bx - ay = d

OpenStudy (amistre64):

you are given a point; and some slope information; use slope point form after using the slope formula to determine the slope

OpenStudy (anonymous):

Okay, I see one formula in my book that says "General Form" Ax + By + C = 0

OpenStudy (amistre64):

yep; that is a general form

OpenStudy (anonymous):

Wow, this one is actually confusing me..darnit..i'm really getting it too

OpenStudy (amistre64):

(-2, -3) -( 2 , 5 ) --------- -4, -8; slope = y/x = -8/-4 = 2

OpenStudy (amistre64):

lol ..... i messed that one up

OpenStudy (anonymous):

Wow, i'm lost..ugh

OpenStudy (amistre64):

y-3=2(x+1)

OpenStudy (anonymous):

Okay, so point-slope formula first

OpenStudy (amistre64):

slope formula first; cant plug in a slope unless you get a slope first

OpenStudy (anonymous):

Okay, just to verify....y=mx+b (point slope comes first)

OpenStudy (anonymous):

Then, the Actual slope-intercept form? y=mx+b?

OpenStudy (anonymous):

Okay, so y-3=2(x+1) is the actual point slope formula....and y=mx+b is the slope-intercept form

OpenStudy (amistre64):

yep; slope formula, then point slope, then slope intercept

OpenStudy (amistre64):

y - 2x +2 +3 y = 2x +5

OpenStudy (anonymous):

I guess that is where i'm getting confused..I am using the (-2, -3) and (2, 5) for this formula right?

OpenStudy (anonymous):

The second part assuming...i'm actually lost here

OpenStudy (anonymous):

Hey, I know you are busy and you have been an enormous help. Thank you so very much. I appreciate all your hard work. Makes so much more sense now.